[Hallc_running] [New Logentry] RC report - 7/20/22

arunts at jlab.org arunts at jlab.org
Wed Jul 20 10:45:02 EDT 2022

Logentry Text:
We collected 32k pions on LH2 in the previous configuration and successfully moved on to the Q^2 = 8.5 GeV^2 data point early during yesterday's day shift. Tests were done by the MCC while we were changing configuration. We continued to have many beam trips after these tests. Data taking was then interrupted by thunderstorms and power supply issues afterwards. We did not get any beam throughout most of yesterday's swing and owl shifts due to these issues and we transitioned to 'maintenance day' this morning. Walter and tech, Heidi and Jixie are working on their respective tasks in the hall which shouldn't take more than a few hours. The hope is that beam returns sometime during swing shift today and we resume production running with stable beam current. 

Let's have a brief RC meeting today at 11:00 AM EST.

Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4017821
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220720/135cbf39/attachment.html>

More information about the Hallc_running mailing list